# An Introduction on Solar Imaging Data Analytic Challenges and Opportunities for Statisticians

### v1

@@ 2.5 SOHO and GONG: 2D Solar Imaging Data
Figure 10 shows the GONG magnetogram data from the NSO website, accessed on May 3, 2024.
@@ 2.6 Existing Data Products and Software
In the space weather literature, the data products are all open-access and available for researchers to download. However, the pre-processing of the raw data requires quite a bit of efforts and domain expertise. In response to this, there has been research works on pre-processed ML-ready data for solar flare predictions. For example, Angryk et al. (2020b) publishes “a comprehensive, multivariate time series (MVTS) dataset extracted from solar photospheric vector magnetograms in Spaceweather HMI Active Region Patch (SHARP) series”. The dataset covers 4,098 MVTS data collections from active regions occurring between May 2010 and December 2018, includes 51 flare-predictive parameters, and integrates over 10,000 flare reports. However, this data product is not on raw solar imaging but summary statistics of solar patches, despite its nice properties of data pre-processing.
@@
Stanford University’s Joint Science Operation Center (JSOC) stores data from SoHO MDI, SDO HMI and AIA, and various other solar instruments. The SunPy-affiliated package DRMS enables the querying of these images (Glogowski et al., 2019; Barnes et al., 2020). All of these individual image products from JSOC are at the same processing level and are supplied in a Flexible Image Transport System (FITS) format that contains only scalar values. The NASA Solar Data Analysis Center’s (SDAC) Virtual Solar Observatory(Hill et al., 2009) (VSO) tool enables data queries from a number of individual data providers.
Another source of data prepared in our recently developed pre-processing pipeline (to be published) is the Atmospheric Imaging Assembly (AIA) (Lemen et al., 2012; Boerner et al., 2012), which provides images of the solar corona in the extreme ultraviolet at relevant wavelengths (9.4nm, 13.1nm, 17.1nm, 19.3nm, 21.1nm, 30.4nm and 33.5nm). These data can enrich the information of each active region with images with emission in the low corona.
@@
In contrast to the SDOML that down-sampled the full-disk images in space and time to keep the dataset size manageable, AARP database preserves the native spatial resolution of SDO/AIA well samples the temporal evolution of solar active regions to match the SDO/HMI HARP database (Dissauer et al., 2023; Leka et al., 2023). The current AARP database includes daily 7 hours samples of 13 minutes of images (centered hourly on “*.48UT” from 15:48 to 21:48 UT) from 2010 June to 2018 December. The database aims to capture both the short-term dynamics (72 s temporal resolution in 13 minutes sample data) and long-term changes (7 hours of evolution per day). The total size of the database is 9 TB. This database is available through the NASA Solar Data Analysis Center (SDAC, https://hpde.io/NASA/NumericalData/SDO/AIA/NWRA/AARP/PT12S).
@@
Another SDO dataset that will be used in the studies mentioned in Section 5 is based on a similar idea as AARP but tailored particular for ML-based flare prediction studies (Jin et al. 2024, in prep). Starting from the existing HMI HARP product, the AIA observations in 8 channels (94, 131, 171, 193, 211, 304, 335, 1600 Å) are processed for the same FOV and coordinate (i.e., CEA: cylindrical equal area). The dataset includes all M-class+ flares from 2010-2018, starting 24 hours before the flare onset until flare peak time with a temporal resolution of 12 minutes. A contract database for ML model training is creased by including all B-class flares occurred in the same active regions of the M-class+ flares. In addition, we derive Differential Emission Measure (DEM) maps 1 hour before each flare that could provide additional information for the ML model. The total size of the dataset is 9 TB.
Besides these aforementioned data products, scientists have been actively working on building ML-ready data products to be made publicly available in future. This is exemplified by the NASA call for proposals “Heliophysics Artificial Intelligence/Machine Learning-Ready Data” (NNH23ZDA001N-HARD, ROSES-23 B.16 HARD) in both 2023 and 2024.
@@ 3 Current Work with Solar Imaging Data
In this section, we describe the current literature progress of the machine learning approaches for handling solar data, including those that work with summary statistics instead of raw solar imaging data. Table 7 we provide a brief summary of the (partial) literature on solar flare predictions, including a partial list of representative papers published in recent years on this topic. In the next few subsections, we will describe in more details the common practices that researchers in space weather have adopted to approach the solar flare predictions problem, focusing mainly on our previous work on solar flare forecasting, as an example of utilizing solar data for prediction models.
@@ 3.1 Feature Engineering Prior to Forecasting
The most popular approach that researchers have adopted when handling solar imaging data is to perform feature engineering (e.g., Jonas et al., 2018), based on either physics or machine learning, prior to forecasting models. This efficiently reduces the dimensionality of the input space, which results in less over-fitting, cutting the data volume significantly and reducing the computational burden for prediction model training. In Chen et al. (2019), an autoencoder network is trained on HMI 3D magnetic field data, and the extracted features from the autoencoder are used to forecast strong solar flares. It is shown that the autoencoder can “reconstruct” the 3D magnetogram with a very small root mean squared error, but the extracted features do not outperform the known list of physics parameters (SHARPs (Bobra et al., 2014b)) in terms of solar flare forecasting. A closer examination of the results shows that the reconstructed imaging data looks much smoother than the original observations. In the solar flare mechanism, the local information around the “polarity inversion line” is very important for flare forecasting (Schrijver, 2007). Therefore, black-box feature engineering may not be the best option to extract important features for strong solar eruptions from massive solar images with only a few hundred strong solar flare events.
@@ 3.2 Prediction Model with Multi-channel Segmented Solar Images
The structured information of the high dimensional multi-way tensor data makes the analysis an intriguing but challenging topic for statisticians and practitioners, especially in the context of physical sciences. In the solar flares prediction problem, the observations come as a time series of 3-way tensor of active regions (see section 2.4) of the Sun that produces flares: 3-D magnetosphere maps from the Helioseismic and Magnetic Imager and multi-channel temperature maps from the Atmospheric Imaging Assembly (Bobra et al., 2014b). Figure 12 shows eight AIA channels and one HMI channel and one Polarity Inversion Line (PIL) for an active region on the Sun that produced an M-class (strong) flare. The data size is for this one-time point, and we have data from 2010 till now at every 12 minutes.
In our recent paper in ICML, Sun et al. (2023b), we expand the Tensor-GP model by integrating a dimensionality reduction technique, called tensor contraction, with a Tensor-GP for a scalar-on-tensor regression task with multi-channel imaging data. We first estimate a latent, reduced-size tensor for each data tensor and then apply a multi-linear Tensor-GP on the latent tensor data for prediction. We introduce an anisotropic total-variation regularization when conducting the tensor contraction to obtain a sparse and smooth latent tensor. We then propose an alternating proximal gradient descent algorithm for estimation. We validate our approach via extensive simulation studies and applying it to the solar flare forecasting problem. More precisely, we consider a multi-channel imaging dataset , where with as the height, width and number of channels, respectively; and . We use to denote the channel of . Our model is specified as
@@ 3.3 Prediction Model with Full Disc Solar Images
Despite the fact that it is known based on heliophysics knowledge, that energetic solar flares are localized in small patches of the Sun, named active regions (ARs, see section 2.4 for details and references), there are continuous efforts among machine learning and heliophysics community on constructing prediction models with full disc solar images, with the hope that the machine can figure out the active regions in an automated fashion. The pros of this approach are that it is not restricted to established knowledge and thus has the potential of identifying solar flares that are not localized in active regions if there exists any. The cons of this approach is that it needs to rely on the limited number of strong solar flares to construct very sparse features from very high-dimensional imaging data, which can result in over-fitting when training machine learning models. However, if adopted properly, the authors believe that the large volume of raw solar imaging data for both quiet time and flaring time can potentially be utilized to train a large black-box machine learning model with competitive performances. Whether or not this model will outperform those models informed by physics knowledge or supplemented by laws of plasma physics remains unknown.
Pandey et al. (2023a) uses CNN for “hourly full-disk line-of-sight magnetogram images and employs a binary prediction mode to forecast M-class flares that may occur within the following 24-h period”. The study shows that the full-disc analysis is aligned with precursors that occur in active regions. It is claimed that the trained model can learn shape and texture-based characteristics, even if in near-the-limb regions, where majority of the previous literature do not consider (e.g. Chen et al., 2019). Furthermore, in a followup work by the same group, Pandey et al. (2023b), focuses on the near-the-limb regions and explored multiple networks for strong flare forecasting.
@@ 4 Looking into the Future: Gaps and Opportunities
In summary, an extensive set of machine learning algorithms have been tried out to show promising results of data-driven solar flare forecasting. Other solar eruptions such as the more severe solar energetic particles have also been tested with data driven approaches (Kasapis et al., 2022; Whitman et al., 2022; Kasapis et al., 2024). However, majority of the works directly convert/prepare the solar data into a classical binary classification problem, ignoring various levels of complications of the solar imaging data and solar flaring properties. We list a few of the major features of the solar imaging and flaring data here, thus pointing out challenges and opportunities for developments in statistical theory, methodology and computational algorithms.
@@ 4.1 Solar Cycle
The Sun’s activity follows approximately 11-year cycles, which is shown in Figure 3, represented by the sunspot numbers. During the peak of a solar cycle, called solar max, energetic solar eruptions are far more frequent than during the valley of a solar cycle, called solar minimum. Qahwaji and Colak (2007) uses the sunspot groups and solar cycle data to forecast strong solar flares. Wang et al. (2020) examines the solar cycle dependency of solar flare forecasting with the LSTM method, showing statistically significant variation when different years of data are chosen for training and testing. This leads to issues when comparing different models in the literature. Furthermore, the high-quality observations, for example, the SDO data, only cover one solar cycle. This makes it hard to learn repeating patterns for solar eruptions from purely data-driven approaches. Statistical models, especially Bayesian statistical methods, will turn out to be efficient in incorporating such information, especially considering operational settings. Furthermore, it is important to know that the solar cycle is not a deterministic quantity (periodicity). It is shown in Figure 3 that we are currently experiencing a much stronger solar cycle than projected (predicted by NASA). Therefore, accounting for the uncertainty in solar cycle forecasting, within the solar eruption forecasting, is also of vital importance. Figure 13 shows the maximum F10.7 cm radio flux (typically three are recorded each day) against the proportion of minutely GOES X-ray fluxes at or above the C-class threshold of .
@@ 4.2 Heterogeneity
The heterogeneity of properties of active regions and flaring mechanisms, on top of the solar cycle dependence, contribute to the difficulties of predicting strong solar flare events. The different active regions have very different “lifetime”, very different sizes, and can erupt drastically different number of solar flares. The determination of the lifetime of active regions is also biased by the fact that direct observations are limited to the time when the region is located in the solar hemisphere facing Earth, and thus, the regions could emerge and/or disappear during the time they are located on “far-side” of the Sun.
Figure 14 shows histograms of active region lifetime and pixelsizes, for active regions from 05/01/2010 till 02/24/2024. Furthermore, Table 8 shows the number of active regions (ARs) corresponding to the specified number (1, 2, 3, 4, 5, and ) of weak (B) and strong (M/X) flare events for each active region recorded in the GOES data set from 05/01/2010 till 06/20/2018. The active region information (i.e., size) is typically used in flare forecasting as a predictor. However, in majority of the works in literature, “personalized” forecasting for each active region as it emerges and fades has not been thoroughly investigated. An earlier paper, Wheatland (2004), proposed a Bayesian method that uses the flaring record of an active region to refine an initial prediction for the occurrence of a big flare during a subsequent period of time. This shows the initial success of solar flare forecasting when taking into account active region evolution.
@@ 4.5 Computational Efficiency
Despite the sample sparsity problem, the raw (most likely multi-channel) imaging data that is observed every few seconds or minutes spanning over the range of a decade or two can be large in volume (e.g., tens of terabytes). The computational efficiency of statistical and machine learning methods plays an important role in the successful implementations of the algorithms on solar imaging data.
@@ 4.6 Uncertainty Quantification
Properly quantifying the uncertainty for space weather monitoring and forecasting is important for risk assessment and decision making (Licata and Mehta, 2022). Probabilistic prediction of solar eruptions should include not only the probability of observing a major eruption in a future time, as given in the majority of the papers in the literature, but also an uncertainty estimate of the probability. Given a major eruption, the uncertainty estimate of the intensity of a solar eruption is also of importance. For example, the peak X-ray flux intensity of a strong solar flare can be an important predictor for a solar proton event (Kasapis et al., 2022).

### v2

@@ 2.5 SOHO and GONG: 2D Solar Imaging Data
Figure 10 shows the GONG magnetogram data from the NSO website, accessed on 3 May 2024.
@@ 2.6 Existing Data Products and Software
In the space weather literature, the data products are open-access and available to researchers to download. However, preprocessing of the raw data requires quite a bit of domain science and machine learning expertise. In response to this, research has been done on pre-processed ML-ready data for solar flare predictions. Angryk et al. (2020b) publishes “a comprehensive, multivariate time series (MVTS) dataset extracted from solar photospheric vector magnetograms in Spaceweather HMI Active Region Patch (SHARP) series.”. The dataset covers 4,098 MVTS data collections from active regions occurring between May 2010 and December 2018, includes 51 flare-predictive parameters, and integrates more than 10,000 flare reports. However, this data product is not based on raw solar imaging but on summary statistics of solar patches, despite its nice properties of data preprocessing.
Table 7 links to commonly used code to download, process, and visualize the solar imaging data products described in the subsequent paragraphs in this section.
@@
Stanford University’s Joint Science Operation Center (JSOC) stores data from SoHO MDI, SDO HMI and AIA, and various other solar instruments. The SunPy-affiliated package DRMS enables querying these images (Glogowski et al., 2019; Barnes et al., 2020). These individual image products from JSOC are at the same processing level and are supplied in a Flexible Image Transport System (FITS) format containing only scalar values. The NASA Solar Data Analysis Center’s (SDAC) Virtual Solar Observatory(Hill et al., 2009) (VSO) tool enables data queries from a number of individual data providers.
@@
In contrast to the SDOML that down-sampled the full-disk images in space and time to keep the dataset size manageable, the AARP database preserves the native spatial resolution of SDO/AIA well samples the temporal evolution of solar active regions to match the SDO/HMI HARP database (Dissauer et al., 2023; Leka et al., 2023). The current AARP database includes daily 7-hour samples of 13 minutes of images (centered hourly on “*.48UT” from 15:48 to 21:48 UT) from June 2010 to December 2018. The database aims to capture both the short-term dynamics (72 seconds’ temporal resolution in 13 minutes’ sample data) and long-term changes (7 hours of evolution per day). The total size of the database is 9 TB. This database is available through the NASA Solar Data Analysis Center (SDAC).
@@
Another SDO dataset that will be used in the studies mentioned in Section 5 is based on a similar idea as AARP but tailored particularly for ML-based flare prediction studies (Jin et al. 2024, in prep). Starting from the existing HMI HARP product, the AIA observations (Lemen et al., 2012; Boerner et al., 2012) in 8 channels (94, 131, 171, 193, 211, 304, 335, 1600 Å) are processed for the same FOV (Field of View) and coordinate (i.e., CEA: cylindrical equal area). The dataset includes all B-class, C-class, and M-class flares from 2010-2024, starting 24 hours before the flare onset until flare peak time with a temporal resolution of 12 minutes. In addition, we derive the Differential Emission Measure (DEM) maps 1 hour before each flare, which could provide additional information for the ML model. The total size of the dataset is 9 TB.
In addition to these aforementioned data products, scientists have been actively working on building ML-ready data products that will be made publicly available in the future. This is exemplified by the NASA call for proposals “Heliophysics Artificial Intelligence / Machine Learning Ready Data” (NNH23ZDA001N-HARD, ROSES-23 B.16 HARD) in 2023 and 2024.
@@ 3 Current Work with Solar Imaging Data
In this section, we describe the current literature on the progress of machine learning approaches for handling solar data, including those that work with summary statistics instead of raw solar imaging data. Table 8 we provide a brief summary of the (partial) literature on solar flare predictions, including a partial list of representative papers published in recent years on this topic. In solar flare predictions literature, the majority of the work adopts standard machine learning models, as given in the table. The predicted quantity is either the binary indicator of a strong or weak flare or the (logarithm of) peak flux intensity of a flare event. The predictors are either summary statistics (e.g., the SHARP parameters) or raw solar imaging data. Cross entropy loss and mean squared error loss are the most commonly adopted loss functions in the solar flare prediction literature. In binary classification, i.e., strong/weak flare prediction, because the samples are highly imbalanced (the number of strong flares is much smaller than the number of weaker flares), researchers have been using the HSS (Heidke skill score) and TSS (true skill score) as the metric for performance evaluation of prediction models (see e.g., Chen et al., 2019). It is still an ongoing endeavor for researchers to obtain fair comparisons of the performances of machine learning models for solar flare forecasting. The difficulty lies primarily in the fact that different researchers/papers process and prepare the training, validation and testing data in their own ways, resulting in a lack of fair comparisons of results despite using the same metrics. For example, splitting data randomly versus splitting by non-overlapping years or splitting by active regions can result in significantly different results, as noted in our previous studies (Wang et al., 2020). Furthermore, due to the scarcity of samples of strong solar flares, there are also non-negligible variations among different sample splits.
In the following few subsections, we will describe in more detail the common practices that researchers in space weather have adopted to approach the solar flare prediction problem, focusing mainly on our previous work on solar flare forecasting as an example of utilizing solar data for prediction models.
@@ 3.1 Feature Engineering Prior to Forecasting
The most popular approach that researchers have adopted when handling solar imaging data is to perform feature engineering (e.g., Jonas et al., 2018), based on either physics or machine learning, prior to forecasting models. This efficiently reduces the dimensionality of the input space, which results in less over-fitting, cutting the data volume significantly and reducing the computational burden for prediction model training. In Stenning et al. (2013), morphological image analysis (Soille, 1999) techniques are adopted for solar images. It is shown that through the morphological image analysis, scientifically meaningful and interpretable numerical features can be extracted from high-throughput solar images for downstream classification and prediction tasks. In Chen et al. (2019), an autoencoder network is trained on HMI 3D magnetic field data, and the extracted features from the autoencoder are used to forecast strong solar flares. It is shown that the autoencoder can “reconstruct” the 3D magnetogram with a very small root mean squared error, but the extracted features do not outperform the known list of physics parameters (SHARPs (Bobra et al., 2014b)) in terms of solar flare forecasting. A closer examination of the results shows that the reconstructed imaging data looks much smoother than the original observations. In the solar flare mechanism, the local information around the “polarity inversion line” is very important for flare forecasting (Schrijver, 2007). Therefore, black-box feature engineering may not be the best option to extract important features for strong solar eruptions from massive solar images with only a few hundred strong solar flare events.
@@ 3.2 Prediction Model with Multi-channel Segmented Solar Images
The structured information of the high dimensional multi-way tensor data makes the analysis an intriguing but challenging topic for statisticians and practitioners, especially in the context of physical sciences. In the solar flares prediction problem, the observations come as a time series of the 3-way tensor of active regions (see section 2.4) of the Sun that produces flares: 3-D magnetosphere maps from the Helioseismic and Magnetic Imager and multi-channel temperature maps from the Atmospheric Imaging Assembly (Bobra et al., 2014b). Figure 12 shows eight AIA channels, one HMI channel, and one Polarity Inversion Line (PIL) for an active region on the Sun that produced an M-class (strong) flare. The data size is for this one-time point, and we have data from 2010 till now every 12 minutes.
In our recent paper in ICML, Sun et al. (2023b), we expand the Tensor-GP model by integrating a dimensionality reduction technique, called tensor contraction, with a Tensor-GP for a scalar-on-tensor regression task with multi-channel imaging data. We first estimate a latent, reduced-size tensor for each data tensor and then apply a multi-linear Tensor-GP on the latent tensor data for prediction. We introduce an anisotropic total-variation regularization when conducting the tensor contraction to obtain a sparse and smooth latent tensor. We then propose an alternating proximal gradient descent algorithm for estimation. We validate our approach via extensive simulation studies and apply it to the solar flare forecasting problem. More precisely, we consider a multi-channel imaging dataset , where the multichannel solar images with as the height, width and number of channels, respectively; and the logarithmic peak flare intensity . We use to denote the channel of . Our model is specified as
@@ 3.3 Prediction Model with Full Disc Solar Images
Despite the fact that it is known based on heliophysics knowledge, that energetic solar flares are localized in small patches of the Sun, named active regions (ARs, see section 2.4 for details and references), there are continuous efforts among machine learning and heliophysics community on constructing prediction models with full disc solar images, with the hope that the machine can figure out the active regions in an automated fashion. The pros of this approach are that it is not restricted to established knowledge and thus has the potential of identifying solar flares that are not localized in active regions if there exist any. The cons of this approach is that it needs to rely on the limited number of strong solar flares to construct very sparse features from very high-dimensional imaging data, which can result in over-fitting when training machine learning models. However, if adopted properly, the authors believe that the large volume of raw solar imaging data for both quiet time and flaring time can potentially be utilized to train a large black-box machine learning model with competitive performances. Whether or not this model will outperform those models informed by physics knowledge or supplemented by laws of plasma physics remains unknown.
Pandey et al. (2023a) uses CNN for “hourly full-disk line-of-sight magnetogram images and employs a binary prediction mode to forecast M-class flares that may occur within the following 24-h period”. The study shows that the full-disc analysis is aligned with precursors that occur in active regions. It is claimed that the trained model can learn shape and texture-based characteristics, even if in near-the-limb regions, where the majority of the previous literature does not consider (e.g. Chen et al., 2019). Furthermore, in a follow-up work by the same group, Pandey et al. (2023b), focuses on the near-the-limb regions and explores multiple networks for strong flare forecasting.
@@ 4 Looking into the Future: Gaps and Opportunities
In summary, an extensive set of machine learning algorithms has been tried out to show promising results of data-driven solar flare forecasting. Other solar eruptions, such as the more severe solar energetic particles, have also been tested with data-driven approaches (Kasapis et al., 2022; Whitman et al., 2022; Kasapis et al., 2024). However, the majority of the works directly convert/prepare the solar data into a classical binary classification problem, ignoring various levels of complications of the solar imaging data and solar flaring properties. We list a few of the major features of the solar imaging and flaring data here, thus pointing out challenges and opportunities for developments in statistical theory, methodology, and computational algorithms.
@@ 4.1 Solar Cycle
The Sun’s activity follows approximately 11-year cycles, which is shown in Figure 3, represented by the sunspot numbers. During the peak of a solar cycle, called solar max, energetic solar eruptions are far more frequent than during the valley of a solar cycle, called solar minimum. Qahwaji and Colak (2007) uses the sunspot groups and solar cycle data to forecast strong solar flares. Wang et al. (2020) examines the solar cycle dependency of solar flare forecasting with the LSTM method, showing statistically significant variation when different years of data are chosen for training and testing. This leads to issues when comparing different models in the literature. Furthermore, the high-quality observations, for example, the SDO data, only cover one solar cycle. This makes it hard to learn repeating patterns for solar eruptions from purely data-driven approaches. Statistical models, especially Bayesian statistical methods, will turn out to be efficient in incorporating such information, especially considering operational settings. Furthermore, it is important to know that the solar cycle is not a deterministic quantity (periodicity). It is shown in Figure 3 that we are currently experiencing a much stronger solar cycle than projected (predicted by NASA). Therefore, accounting for the uncertainty in solar cycle forecasting, within the solar eruption forecasting, is also of vital importance. Figure 13 shows the maximum F10.7 cm radio flux (typically three are recorded each day) against the proportion of minutely GOES X-ray fluxes at or above the C-class threshold of . The data ranges from 2004-10-28 to 2024-01-25.
@@ 4.2 Heterogeneity
The heterogeneity of properties of active regions and flaring mechanisms, on top of the solar cycle dependence, contribute to the difficulties of predicting strong solar flare events. The different active regions have very different “lifetimes”, very different sizes, and can erupt drastically different numbers of solar flares. The determination of the lifetime of active regions is also biased by the fact that direct observations are limited to the time when the region is located in the solar hemisphere facing Earth, and thus, the regions could emerge and/or disappear during the time they are located on “farside” of the Sun.
Figure 14 shows histograms of active region lifetime and pixelsizes for active regions from 05/01/2010 till 02/24/2024. Furthermore, Table 9 shows the number of active regions (ARs) corresponding to the specified number (1, 2, 3, 4, 5, and ) of weak (B) and strong (M/X) flare events for each active region recorded in the GOES data set from 05/01/2010 till 06/20/2018. The active region information (i.e., size) is typically used as a predictor in flare forecasting. However, in a majority of the works in literature, “personalized” forecasting for each active region as it emerges and fades has not been thoroughly investigated. An earlier paper, Wheatland (2004), proposed a Bayesian method that uses the flaring record of an active region to refine an initial prediction for the occurrence of a big flare during a subsequent period of time. This shows the initial success of solar flare forecasting when taking into account active region evolution.
@@ 4.5 Computational Efficiency
Despite the sample sparsity problem, the raw (most likely multi-channel) imaging data that is observed every few seconds or minutes spanning over the range of a decade or two can be large in volume (e.g., tens of terabytes). The computational efficiency of statistical and machine learning methods plays an important role in the successful implementations of the algorithms on solar imaging data. As it currently stands in literature, to make models computationally feasible to fit, data with reduced temporal and/or spatial resolution (or even summary statistics, such as the SHARP parameters or extracted features, instead of raw multichannel solar imaging data) are adopted for training/validating/testing purposes in machine learning models. Lai et al. (2021) propose a divide-and-conquer method (with generalized fiducial inference) capable of handling massive datasets and providing uncertainty estimates. They analyzed the solar flare brightness with the proposed approach using the SDO/AIA data from Schuh et al. (2013).
@@ 4.6 Uncertainty Quantification
Properly quantifying the uncertainty for space weather monitoring and forecasting is essential for risk assessment and decision making (Licata and Mehta, 2022). Probabilistic prediction of solar eruptions should include not only the probability of observing a major eruption in a future time, as given in the majority of the papers in the literature but also an uncertainty estimate of the probability. Given a major eruption, the uncertainty estimate of the intensity of a solar eruption is also of importance. For example, the peak X-ray flux intensity of a strong solar flare can be an important predictor for a solar proton event (Kasapis et al., 2022).